SpaceRef

SpaceRef


SMART-1 Status: Ongoing Lunar Operations

Status Report From: European Space Agency
Posted: Thursday, February 16, 2006

image

24 Jan 2006 16:36 Report for period 19 December 2005 to 15 January 2006

Smart-1 operation has been nominal during the Christmas period. The push-broom operations were terminated just before Christmas giving the pace for resuming routine payload operations. The only noticeable event was a glitch in the star tracker software that did not have any impact on operations.

The Integral/XMM transponder has been used to carry out tone ranging radio frequency compatibility tests aiming at two Smart-1 tracking campaigns for the cross support with the Chinese CLTC and the Indian ISRO during 2006.

Future Activities

Future activities are focused on the following: End of mission preparation and impact on the Moon selection Finalisation of the tone ranging compatibility tests and validation using ESA stations

Orbital Information

SMART-1 OD396 Close to Apolune 1790
Epoch (UTC) 2006/01/16 10:49:26.1

Elements WRT Moon and its equator of date

Pericentre Distance (km) 2380.022541
Apocentre Distance (km) 4456.635279
Semi Major Axis (km) 3418.328910
Eccentricity 0.303747
Inclination (°) 90.116350
Ascending Node (°) 238.470099
Argument of Pericentre (°) 261.884063
True Anomaly (°) 179.999996
Osculating Orbital Period (h) 4.981682

The changes since apolune1756 as follows: semi-major axis +0.5 km perilune height +2.8 km apolune height -1.8 km orbital period +0.1 min

AOCS

Startracker Performance The plot below shows the startracker CCD temperatures for the period covered by this report. Table As expected CHU2 has suffered from high temperatures during the pushbroom operations that ended on 20 December 2005.

Temperature drop in startracker unit 2 as push-broom operations finish

Electric Propulsion, Power and Thermal

Power The Power Subsystem has performed very well during the reporting period. The last eclipse was a moon penumbra from 11:48:28 to 11:46:16 UTC on the 24 December 2005, with the batteries performing nominally.

The eclipse season is scheduled to start again on the 15 March 2006 at 00:15:15 UTC, with an earth penumbra of a duration of 01:17:41 hours.

Electric Propulsion EP remained off after ending the re-boost phase.

Thermal The Thermal Subsystem has performed very well during the reporting period. The last Push-Broom manoeuvre was on the 19 December 2005 from 17:41:07 to 18:27:07, with the thermal mode High to behave as Low to keep the temperatures of the +Y/-Y panels lower to avoid as much as possible start tracker invalidations.

The next Push-Broom period is scheduled to start again around April 2006, when the sun incidence on the +Y/-Y panels during Push-Broom is bellow 30 degrees. This Push-Broom period is planned to be about 3 months long, as was the last one.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter